OpenStudy (anonymous):

3(x - 6) + 2 = 4(x + 2) - 21 x =

OpenStudy (nurali):

3(x - 6) + 2 = 4(x + 2) - 21 3x-18+2=4x+8-21 3x-16=4x-13 3x-4x=-13+16 -x=3 x=-3

OpenStudy (anonymous):

love u

OpenStudy (nurali):

No problem.